# Equivalence Relations on Vertex Operator Algebras, II: Witt Equivalence and Orbifolds

### v1

@@ Abstract.
Inspired by results in the theory of even lattices, and also the theory of tensor categories, we say that two strongly rational vertex operator algebras are Witt equivalent if their central charges agree and if their representation categories are Witt equivalent as modular tensor categories. Two RCFTs are said to be Witt equivalent if their central charges agree and if their associated 2+1d topological field theories can be separated by a topological surface. We argue that Witt equivalence is necessary for two theories to be related by topological manipulations. We conjecture that it is also sufficient, and give proofs in various special cases.
We relate this circle of ideas to the problem of classifying RCFTs, and to lore concerning deformation classes of quantum field theories. We use the notion of Witt equivalence to argue, assuming the conjectural classification of unitary, $c=1$ RCFTs, that all of the finite global symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible. Finally, we sketch a “quantum Galois theory” for chiral CFTs, which generalizes prior mathematical literature by incorporating non-invertible symmetries; we illustrate this non-invertible Galois theory in the context of the monster CFT, for which we produce a Fibonacci symmetry.
@@ 1.1. For Mathematicians
Witt Equiv. (Def. 3.2)Orbifold Equiv. (Def. 4.3)Strong Witt Equiv. (Def. 3.15)Inner Orb. Equiv. (Def. 4.18)Bulk Genus (Def. 2.14)Hyperbolic Genus (Def. 2.15)Prop. 5.12VOA ConceptsConj. 5.6Thm. 5.1Conj. 5.11Prop. 5.19Prop. 3.16Rem. 4.20Prop. 3.19Rem. 3.5Thm. 5.17[MR24, Thm. LABEL:MR1cor:hypbulk]GeneralizeLattice ConceptsWitt Equiv. (Def. 2.3)Rational Equiv. (Def. 2.4)Genus (Def. 2.1)
@@
for some spherical fusion category $\mathcal{F}$ (see 2.10). We denote the Witt class of a modular tensor category $\mathcal{C}$ by $[\mathcal{C}]$. Without motivating this definition here much further, we note that it reduces exactly to Witt equivalence of metric groups, 2.2, if one considers the (pseudo-unitary) pointed modular tensor categories $\mathcal{C}(D)$, which are in bijection with metric groups $D$ [JS93, EGNO15].
@@
Based on the correspondence between conformal extensions of strongly rational vertex operator algebras and condensable algebras in modular tensor categories [KO02, HKL15, CMSY24] (see 2.18), we can reformulate Witt equivalence as follows, also assuming for simplicity that the vertex operator algebras are positive, i.e. that all irreducible modules except for the adjoint module are positively graded. This is depicted in Figure 10.
@@
We point out (see 4.6) that being related by a common strongly rational conformal subalgebra is likely not an equivalence relation, forcing us to explicitly build in the transitivity into our definition. The potential counterexample follows from the conjectural classification of the positive, strongly rational vertex operator algebras of central charge $c=1$ (see 3.26) and is ultimately a consequence of the famous ADE-classification of the finite subgroups of $\mathrm{SL}(2,\mathbb{C})$, which correspond to Du Val singularities.
There is also the “classical” (or strong) version of orbifold equivalence (see 4.18 and 5.18). Two strongly rational vertex operator algebras $V$ and $V^{\prime}$ are inner orbifold equivalent if they share a strongly rational conformal subalgebra $W$ such that the concrete realization of $W$ inside $V$ (and similarly for $V^{\prime}$) can be chosen such that $W$ and $V$ have the same Heisenberg commutant. Here, it actually suffices to demand that $V$ and $V^{\prime}$ be connected “in one step” in order to obtain an equivalence relation. We have chosen the terminology “inner” because, in some special cases, inner orbifold equivalence reduces to a (group-like) orbifold construction under inner automorphisms.
@@
We expect the forward direction in 5.6 to be much harder to establish. Indeed, we argue in 5.16 that, as a very narrow special case, this would likely imply the uniqueness of the moonshine module [FLM88], 2.20.
@@
5.14. Assuming 2.20 on the uniqueness of the moonshine module, 5.6 is true for the infinitely many positive, strongly rational vertex operator algebras belonging to the Witt classes $([\operatorname{Vect}],c)$ with $c\leq 24$.
@@
Finally, in Section 7, we discuss the relationship between Witt equivalence and generalized global symmetries, through a concept called symmetry-subalgebra duality. While largely written at a physics level of rigor, we expect that much of this can be made into mathematics. In particular, we argue that all the finite symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible [CLR24] (see 7.8), and we also find a Fibonacci symmetry of the moonshine module $V^{\natural}$ (see 7.7, and also [FEH24], which appeared independently during the work on this manuscript.).
@@ 1.2. For Physicists
In the context of 2+1d topological quantum field theories (TQFTs), it was recently rigorously established that two theories can be related by a kind of generalized orbifold procedure [MR22] if and only if they can be separated by a topological interface [Mul24]. Whether two TQFTs admit a topological interface between them is in turn sharply characterized within the framework of tensor categories via the notion of Witt equivalence [DNO13, DMNO13], see 2.10.
@@
    One works with a spacetime manifold of the form $\Sigma\times[-1,1]$, with $\Sigma$ a two-dimensional surface. The region $\Sigma\times[0,1]$ is in the phase defined by the TQFT $\operatorname{Rep}(W)$, and $\Sigma\times[-1,0]$ is in the phase defined by $\operatorname{Rep}(V)$.
@@
5.1. The triple $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ associated with an RCFT ${{}_{V}}\mathcal{H}_{W}$ is invariant under topological manipulations. Thus, if $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ is not equal to $(c(V^{\prime}),c(W^{\prime}),\mathsf{W}(\mathcal{H}^{\prime}))$, then the RCFTs ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ cannot be related by gauging a finite subpart of a generalized global symmetry, or separated by a topological line interface of finite quantum dimension.
The power of 5.1 rests on the fact that the triple $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ is a very coarse shadow of a full RCFT, and correspondingly it is far easier to work with. Furthermore, category-theoretic techniques from 2+1d TQFT can be used to compute $\mathsf{W}(\mathcal{H})$, and matching this invariant is generally more effective than trying to discover whether there exists a sequence of orbifolds, or a topological line interface, which interpolates between two given 1+1d QFTs.
@@
    The topological line interface $I$ inflates into a pair of two-dimensional topological interfaces $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, where $\mathcal{I}_{1}$ interpolates between $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$, and $\mathcal{I}_{2}$ interpolates between $\operatorname{Rep}(W)$ and $\operatorname{Rep}(W^{\prime})$.
@@
It is natural to wonder whether $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ defines a complete invariant. In other words, is it guaranteed that ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ can be related by topological manipulations if their invariants $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ and $(c(V^{\prime}),c(W^{\prime}),\mathsf{W}(\mathcal{H}^{\prime}))$ match?
@@
5.8. If two RCFTs ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ have matching invariants $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ and $(c(V^{\prime}),c(W^{\prime}),\mathsf{W}(\mathcal{H}^{\prime}))$, respectively, then one can be obtained from the other via topological manipulations. More succinctly, the triple $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ defines a complete invariant.
@@
First of all, we show in 5.15, using the results of [MS23], that 5.8 is true in the case that $c(V),c(W)\leq 24$ and $\mathsf{W}(\mathcal{H})=[\operatorname{Vect}]$, if one assumes the widely believed conjecture concerning the uniqueness of the moonshine module, 2.20. In particular, for sufficiently low central charges, 5.15 is true for holomorphically factorizing CFTs, and more generally any RCFT whose, e.g., left-moving chiral algebra is attached to a bulk 2+1d TQFT that admits a topological boundary condition.
@@
Having sketched our results, we now explain why one should care about them. One source of motivation comes from the problem of classifying RCFTs. Orbifolding [DHVW85, DHVW86] is one of the two main procedures by which one can produce new CFTs from known ones (the other being the method of taking cosets [GKO85, GKO86]), and it is natural to ask how effective this procedure is at moving one around theory space. Indeed, throughout the decades, there have been various proposals for how these two tools, orbifolding and cosets, can be used in combination to generate all RCFTs from some smaller set of seed theories. We claim that, on the power of one additional mild assumption (called weak reconstruction, 2.21), one is led to the following “structure theorem” on RCFTs.
6.1. Assuming 5.8 and weak reconstruction, 2.21, any RCFT can be obtained from $\smash{\mathsf{E}_{8,1}^{\otimes n}\otimes\overline{\mathsf{E}}_{8,1}^{\otimes m}}$ for some $n,m\in\mathbb{Z}_{\geq 0}$ by performing topological manipulations and then quotienting decoupled degrees of freedom.
@@
Finally, it is natural to extend the notion of Witt equivalence of 3d TQFTs to 2d RCFTs by defining two theories ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ to be Witt equivalent if their invariants $(c(V),c(W),\mathsf{W}(\mathcal{H}))$ and $(c(V^{\prime}),c(W^{\prime}),\mathsf{W}(\mathcal{H}^{\prime}))$ are equal. A natural question then is to understand the space of all theories that are Witt equivalent to a fixed RCFT ${{}_{V}}\mathcal{H}_{W}$, i.e. to understand the Witt class of ${{}_{V}}\mathcal{H}_{W}$. We explain in 7.4 how this problem is closely related to the problem of classifying the global symmetries of ${{}_{V}}\mathcal{H}_{W}$. We emphasize that this perspective has practical implications by showing that it leads to a proof of the following result.
@@
All Lie algebras and vertex (operator) algebras will be over the base field $\mathbb{C}$. All categories will be enriched over $\operatorname{Vect}=\operatorname{Vect}_{\mathbb{C}}$.
@@ Acknowledgments
We thank Yichul Choi, Aaron Hofer, Gerald Höhn, Hannes Knötzele, Zohar Komargodski, Robert McRae, Yuto Moriwaki, Nils Scheithauer, Sahand Seifnashri, Shu-Heng Shao, Yifan Wang and Xiao-Gang Wen for helpful discussions.
@@ 2.1. Lattices and Metric Groups
We discuss this rather simple fact because this transitivity property will fail for the corresponding notion for strongly rational vertex operator algebras (see 4.3, Figure 11 and 4.6).
By definition, for even lattices $L$ and $M$, being isomorphic (or integrally equivalent) implies rational equivalence. Also, it follows from the Hasse principle that $L$ and $M$ are rationally equivalent if and only if all the local quadratic spaces $L\otimes\mathbb{Q}_{p}$ and $M\otimes\mathbb{Q}_{p}$ for $p$ a prime including $\infty$ are isomorphic. In particular, lattices in the same genus are rationally equivalent.

### v2

@@ Abstract.
Inspired by results in the theory of even lattices and tensor categories, we say that two strongly rational vertex operator algebras are Witt equivalent if their central charges agree and if their modular tensor categories are Witt equivalent. Two RCFTs are said to be Witt equivalent if their central charges agree and if their associated 2+1d topological field theories can be separated by a topological surface. We argue that Witt equivalence is necessary for two theories to be related by topological manipulations. We conjecture that it is also sufficient, and give proofs in various special cases.
We relate this circle of ideas to the problem of classifying RCFTs, and to lore concerning deformation classes of quantum field theories. We use the notion of Witt equivalence to argue, assuming the conjectural classification of unitary, $c=1$ RCFTs, that all of the finite symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible. We also sketch a “quantum Galois theory” for chiral CFTs, which generalizes prior mathematical literature by incorporating non-invertible symmetries; we illustrate this non-invertible Galois theory in the context of the monster CFT, for which we produce two Fibonacci lines. Finally, we discuss $p$-neighborhood of vertex operator algebras, which is a special topological manipulation related to $\mathbb{Z}_{p}$-orbifolding.
@@ 1.1. For Mathematicians
Witt Equiv. (Def. 3.2)Orbifold Equiv. (Def. 4.3)Strong Witt Equiv. (Def. 3.15)Inner Orb. Equiv. (Def. 4.19)Bulk Genus (Def. 2.15)Hyperbolic Genus (Def. 2.16)Prop. 5.12VOA ConceptsConj. 5.6Thm. 5.1Conj. 5.11Prop. 5.19Prop. 3.16Rem. 4.21Prop. 3.19Rem. 3.5Thm. 5.17[MR24, Thm. LABEL:MR1cor:hypbulk]GeneralizeLattice ConceptsWitt Equiv. (Def. 2.3)Rational Equiv. (Def. 2.4)Genus (Def. 2.1)
@@
for some spherical fusion category $\mathcal{F}$ (see 2.11). We denote the Witt class of a modular tensor category $\mathcal{C}$ by $[\mathcal{C}]$. Without motivating this definition here much further, we note that it reduces exactly to Witt equivalence of metric groups, 2.2, if one considers the (pseudo-unitary) pointed modular tensor categories $\mathcal{C}(D)$, which are in bijection with metric groups $D$ [JS93, EGNO15].
@@
Based on the correspondence between conformal extensions of strongly rational vertex operator algebras and condensable algebras in modular tensor categories [KO02, HKL15, CMSY24] (see 2.19), we can reformulate Witt equivalence as follows, also assuming for simplicity that the vertex operator algebras are positive, i.e. that all irreducible modules except for the adjoint module are positively graded. This is depicted in Figure 11.
@@
We point out (see 4.7) that being related by a common strongly rational conformal subalgebra is likely not an equivalence relation, forcing us to explicitly build in the transitivity into our definition. The potential counterexample follows from the conjectural classification of the positive, strongly rational vertex operator algebras of central charge $c=1$ (see 3.26) and is ultimately a consequence of the famous ADE-classification of the finite subgroups of $\mathrm{SL}(2,\mathbb{C})$, which correspond to Du Val singularities.
There is also the “classical” (or strong) version of orbifold equivalence (see 4.19 and 5.18). Two strongly rational vertex operator algebras $V$ and $V^{\prime}$ are inner orbifold equivalent if they share a strongly rational conformal subalgebra $W$ such that the concrete realization of $W$ inside $V$ (and similarly for $V^{\prime}$) can be chosen such that $W$ and $V$ have the same Heisenberg commutant. Here, it actually suffices to demand that $V$ and $V^{\prime}$ be connected “in one step” in order to obtain an equivalence relation. We have chosen the terminology “inner” because, in some special cases, inner orbifold equivalence reduces to a (group-like) orbifold construction under inner automorphisms.
@@
We expect the forward direction in 5.6 to be much harder to establish. Indeed, we argue in 5.16 that, as a very narrow special case, this would likely imply the uniqueness of the moonshine module [FLM88], 2.21.
@@
5.14. Assuming 2.21 on the uniqueness of the moonshine module, 5.6 is true for the infinitely many positive, strongly rational vertex operator algebras belonging to the Witt classes $([\operatorname{Vect}],c)$ with $c\leq 24$.
@@
In Section 7, we discuss the relationship between Witt equivalence and generalized global symmetries, through a concept called symmetry-subalgebra duality. While largely written at a physics level of rigor, we expect that much of this can be made into mathematics. In particular, we argue that all the finite symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible [CLR24] (see 7.8), and we also find a Fibonacci symmetry of the moonshine module $V^{\natural}$ (see 7.7, and also [FEH24], which appeared independently during the work on this manuscript).
Finally, in Section 8 we study $p$-neighborhood of vertex operator algebras. This is a specialization of orbifold equivalence where we allow two strongly rational vertex operator algebras $V$ and $V^{\prime}$ to be related by one common strongly rational conformal subalgebra $W$ of index $p$ for a prime $p$ (see 8.1).
This definition is inspired by the notion of neighborhood for lattices [Kne57]. For even lattices, the connected component of the $p$-neighborhood graph for a fixed prime $p$ is usually identical to the genus (see Section 2.1).
However, for strongly rational vertex operator algebras the picture is less clear. For simplicity, we restrict to the case of holomorphic vertex operator algebras. Naively, based on the dichotomy between orbifold and Witt equivalence and their inner versions (see Figure 1), one would expect that iterated $p$-neighborhood corresponds to the bulk genus and iterated inner $p$-neighborhood to the hyperbolic genus. However, the $p$-neighborhood graph of a bulk genus of holomorphic vertex operator algebras is usually disconnected (see 8.17 and [Mon98, HM23]), while one might suspect that it becomes connected if one is allowed to mix the primes $p$ (see 8.2).
The picture becomes clearer in the inner case. In 8.14 we essentially propose that, except for finitely many primes, the inner $p$-neighborhood graph of a hyperbolic genus is connected for a fixed prime $p$. We are able to prove a slightly weaker statement in 8.16.
In Section 8.5, we provide empirical evidence for many of the formulated conjectures pertaining to $p$-neighborhood.
@@ 1.2. For Physicists
In the context of 2+1d topological quantum field theories (TQFTs), it was recently rigorously established that two theories can be related by a kind of generalized orbifold procedure [MR22] if and only if they can be separated by a topological interface [Mul24]. Whether two TQFTs admit a topological interface between them is in turn sharply characterized within the framework of tensor categories via the notion of Witt equivalence [DNO13, DMNO13], see 2.11.
@@
    One works with a spacetime manifold of the form $\Sigma\times[-1,1]$, with $\Sigma$ a $2$-dimensional surface. The region $\Sigma\times[0,1]$ is in the phase defined by the TQFT $\operatorname{Rep}(W)$, and $\Sigma\times[-1,0]$ is in the phase defined by $\operatorname{Rep}(V)$.
@@
5.1. The triple $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ associated with an RCFT ${{}_{V}}\mathcal{H}_{W}$ is invariant under topological manipulations. Thus, if $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ is not equal to $(\mathsf{W}(\mathcal{H}^{\prime}),c(V^{\prime}),c(W^{\prime}))$, then the RCFTs ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ cannot be related by gauging a finite subpart of a generalized global symmetry, or separated by a topological line interface of finite quantum dimension.
The power of 5.1 rests on the fact that the triple $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ is a very coarse shadow of a full RCFT, and correspondingly it is far easier to work with. Furthermore, category-theoretic techniques from 2+1d TQFT can be used to compute $\mathsf{W}(\mathcal{H})$, and matching this invariant is generally more effective than trying to discover whether there exists a sequence of orbifolds, or a topological line interface, which interpolates between two given 1+1d QFTs.
@@
    The topological line interface $I$ inflates into a pair of $2$-dimensional topological interfaces $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, where $\mathcal{I}_{1}$ interpolates between $\operatorname{Rep}(V)$ and $\operatorname{Rep}(V^{\prime})$, and $\mathcal{I}_{2}$ interpolates between $\operatorname{Rep}(W)$ and $\operatorname{Rep}(W^{\prime})$.
@@
It is natural to wonder whether $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ defines a complete invariant. In other words, is it guaranteed that ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ can be related by topological manipulations if their invariants $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ and $(\mathsf{W}(\mathcal{H}^{\prime}),c(V^{\prime}),c(W^{\prime}))$ match?
@@
5.8. If two RCFTs ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ have matching invariants $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ and $(\mathsf{W}(\mathcal{H}^{\prime}),c(V^{\prime}),c(W^{\prime}))$, respectively, then one can be obtained from the other via topological manipulations. More succinctly, the triple $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ defines a complete invariant.
@@
First of all, we show in 5.15, using the results of [MS23], that 5.8 is true in the case that $c(V),c(W)\leq 24$ and $\mathsf{W}(\mathcal{H})=[\operatorname{Vect}]$, if one assumes the widely believed conjecture concerning the uniqueness of the moonshine module, 2.21. In particular, for sufficiently low central charges, 5.15 is true for holomorphically factorizing CFTs, and more generally any RCFT whose, e.g., left-moving chiral algebra is attached to a bulk 2+1d TQFT that admits a topological boundary condition.
@@
Having sketched our results, we now explain why one should care about them. One source of motivation comes from the problem of classifying RCFTs. Orbifolding [DHVW85, DHVW86] is one of the two main procedures by which one can produce new CFTs from known ones (the other being the method of taking cosets [GKO85, GKO86]), and it is natural to ask how effective this procedure is at moving one around theory space. Indeed, throughout the decades, there have been various proposals for how these two tools, orbifolding and cosets, can be used in combination to generate all RCFTs from some smaller set of seed theories. We claim that, on the power of one additional mild assumption (called weak reconstruction, 2.22), one is led to the following “structure theorem” on RCFTs.
6.1. Assuming 5.8 and weak reconstruction, 2.22, any RCFT can be obtained from $\smash{\mathsf{E}_{8,1}^{\otimes n}\otimes\overline{\mathsf{E}}_{8,1}^{\otimes m}}$ for some $n,m\in\mathbb{Z}_{\geq 0}$ by performing topological manipulations and then quotienting decoupled degrees of freedom.
@@
Finally, it is natural to extend the notion of Witt equivalence of 3d TQFTs to 2d RCFTs by defining two theories ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ to be Witt equivalent if their invariants $(\mathsf{W}(\mathcal{H}),c(V),c(W))$ and $(\mathsf{W}(\mathcal{H}^{\prime}),c(V^{\prime}),c(W^{\prime}))$ are equal. A natural question then is to understand the space of all theories that are Witt equivalent to a fixed RCFT ${{}_{V}}\mathcal{H}_{W}$, i.e. to understand the Witt class of ${{}_{V}}\mathcal{H}_{W}$. We explain in 7.4 how this problem is closely related to the problem of classifying the global symmetries of ${{}_{V}}\mathcal{H}_{W}$. We emphasize that this perspective has practical implications by showing that it leads to a proof of the following result.
@@
All Lie algebras and vertex (operator) algebras will be over the base field $\mathbb{C}$. All categories will be enriched over $\operatorname{Vect}=\operatorname{Vect}_{\mathbb{C}}$. Usually, $q$ denotes a formal variable or the complex variable $q=e^{2\pi i\tau}$ for $\tau\in\mathbb{H}$ with the complex upper half-plane $\mathbb{H}\coloneqq\{z\in\mathbb{C}\mid\operatorname{Im}(z)>0\}$.
@@ Acknowledgments
We thank Yichul Choi, Thomas Creutzig, Aaron Hofer, Gerald Höhn, Hannes Knötzele, Zohar Komargodski, Robert McRae, Yuto Moriwaki, Nils Scheithauer, Sahand Seifnashri, Shu-Heng Shao, Yifan Wang, Xiao-Gang Wen, Harshit Yadav and Hiroshi Yamauchi for helpful discussions.
@@ 2.1. Lattices and Metric Groups
We discuss this rather simple fact because this transitivity property will fail for the corresponding notion for strongly rational vertex operator algebras (see 4.3, Figure 12 and 4.7).
By definition, for even lattices $L$ and $M$, being isomorphic (or integrally equivalent) and being $p$-neighbors (see below) each imply implies rational equivalence. Also, it follows from the Hasse principle that $L$ and $M$ are rationally equivalent if and only if all the local quadratic spaces $L\otimes\mathbb{Q}_{p}$ and $M\otimes\mathbb{Q}_{p}$ for $p$ a prime including $\infty$ are isomorphic. In particular, lattices in the same genus are rationally equivalent.
@@
Finally, we also recall the notion of lattice neighborhood [Kne57].
Definition 2.7 (Neighborhood).
For a positive integer $n\in\mathbb{Z}_{>0}$, two even lattices $L$ and $M$ (or isomorphism classes thereof) are called $n$-neighbors, necessarily of the same discriminant $d(L)=d(M)$, if they contain a common full-rank sublattice $K$ of index $n$ in both of them.
That is, $K^{\prime}/K$ has two isotropic subgroups isomorphic to $\mathbb{Z}_{n}$ and the corresponding extensions of $K$ are (isomorphic to) $L$ and $M$. The definition is depicted in Figure 7.
[…]
Figure 7. Neighborhood of even lattices $L$ and $M$.
The lattice genus of $L$, 2.1, is closely related to the connected component of $L$ in the lattice $p$-neighborhood graph when $p$ is a prime not dividing $d(L)$. For instance, if $L$ and $M$ are even lattices that are iterated $p$-neighbors for some prime $p$ with $p\nmid d(L)$, then $L$ and $M$ belong to the same genus.
Conversely, suppose that $L$ and $M$ are even lattices of rank at least $3$ belonging to the same genus, and further assume that this genus consists of only one spinor genus. [^7] In general, the genus of a lattice decomposes into $2^{k}$ many spinor genera for some $k\in\mathbb{Z}_{\geq 0}$. Often, $k=0$, i.e. genus and spinor genus coincide. See, e.g., [O’M71]. Then $L$ and $M$ are iterated $p$-neighbors of one another for any odd prime $p$ with $p\nmid d(L)$; see, e.g., [SP91] and the references cited therein (or [Sch22, Che22]). More precisely, $p$-neighbors of $L$ for a fixed prime $p$ cover (at least) the whole spinor genus of $L$, while different spinor genera in the genus of $L$ may be connected by $p$-neighborhood for some specially selected primes $p$ [BH83, SP91]. We summarize this as follows: considering two even lattices $L$ and $M$ of rank at least 3, if they belong to the same genus, then are they iterated $p$-neighbors of one another for possibly different primes $p$.
We shall consider the generalization of $n$-neighborhood to vertex operator algebras in Section 8, but it seems not to admit such a clear twofold pattern as in Figure 1.